OpenStudy (anonymous):

how would I solve this equation for x? 0=3.9657(0.9982^x)ln0.9982

OpenStudy (anonymous):

\[0=3.9657(0.9982^x)\ln0.9982\]

OpenStudy (welshfella):

first look up ln 0.9982 on your calculator and multiply what you get by 3.9657

OpenStudy (anonymous):

For ln 0.9982, I got -0.0018 and when I multiplied it by 3.9657 I got -0.007145. Would it then be: 0=(0.9982^x)

OpenStudy (ybarrap):

$$ 0=3.9657\left (0.9982^x\right )\ln {0.9982}\\ {0\over 3.9657 \times \ln {0.9982}}={\cancel{3.9657}\left (0.9982^x\right )\cancel{\ln {0.9982}}\over \cancel{3.9657} \times \cancel{\ln {0.9982}}}=0.9982^x\\ \sqrt[x]{0}=\sqrt[x]{0.9982^x}=0.9982\\ 0\ne0.9982 $$ Hence, there are no solutions since 0 can not equal 0.9982.